HA1 – R12

Exercise 5. Charging for storage and handling

Please use table 12.5 on p. 288 of the Richards book for this exercise without any changes given in exercise 4. Assume that the cost of one pallet storage is $10 per week, inbound (intake) handling is $2.50 per pallet and the cost of outbound handling (dispatch) is $4.00 per pallet. All overhead is applied at a rate of 20% to both storage and handling. Assume the week consists of seven days, Monday through Sunday. In a closed-book contract where charges are billed on a weekly basis and include 15% profit applied to all costs, the 3PL uses minimum daily average pallet storage requirement of 100 pallets and minimum weekly inbound and outbound handling requirements of 350 pallets each to base the charges on if the actual daily average usage numbers are below the minima. The charges on the bill are itemized as storage, inbound handling, and outbound handling.

16. What will the customer be billed for storage for that week if the parties use the average daily pallet method (scenario 2)?

17. What is the agreed contract charge per storage of one pallet per week?

18. What will the customer be billed for inbound handling for that week?

19. What is the total charge for that week?

20. In an open book or cost-plus contract, a total profit may be the last item on the bill after the cost subtotal. What will this profit line item be under these conditions for that week?

21. Bonus (extra credit) question. Answer question 16 under the modified terms of exercise 4. 22. Bonus (extra credit) question. Answer question 19 under the modified terms of exercise 4.
Answer #1

Please find below summary of the data given in table 12.5.

Data already given are highlighted in Grey. Data computed are highlighted in Green.

# Pallets left at day end (Storage) are calculated as = #pallets (at beginning of the day) + Intake - Dispatch

For example,

  1. Monday :  #Pallets left at day end (Storage) = 100+100-25 = 175 pallets
  2. Tuesday :  #Pallets left at day end (Storage) = 175+100-75 = 200 pallets


Following are the results:

  • Total Pallets Intake = 300 pallets
  • Total Pallets Dispatch = 330 pallets
  • Average Daily Storage of pallets = 845/7 = 120.714 pallets.

Now, we will calculate the billing amount for these 3 activities.

Please find above table for net bill charge for 3 activities. Cells highlighted in Grey are for the data given in the problem statement. Cells highlighted in Green are computed results.

For example,
Inbound handling: Cost per pallet is $2.5, Overhead charges are 20% and Profit margin is 15%,
Hence, cost including overhead charges would be = 2.5*(1+20%) = 2.5*120% = 3
Final bill amount for 3PL (Including profit) = 3*(1+15%) = 3*115% = $3.45
Likewise, we can calculate final bill amount for remaining 2 activities.

Following are the net bill amounts:

  • Inbound handling: $3.45 per pallet
  • Outbound handling: $ 5.52 per pallet
  • Storage : $ 13.80 per pallet

We will be using above calculated data for answering the questions.

Answer for Question 16:

If scenario 2 i.e. average daily storage is used for billing purpose, No of pallets for billing purpose = 100 pallets.

Bill amount = No of pallets * Bill charge for storage = 100 * 13.8 = $ 1380.

Answer for Question 17:

Agreed contract charge for Storage = $13.8. Please refer to above image for calculation & methodology.

Answer for Question 18:

Inbound Handling charges = No of pallets * Bill charge for inbound handling  

Bill Charge for inbound handling = $3.45 as per calculation done above.

Since No of pallets intake for the week is 300 which is less than minimum contractual agreement of 350 pallets, we will consider 350 pallets for the billing purpose.

Hence, Inbound Handling charges = No of pallets * Bill charge for storage = 350 * 3.45 = $1207.50

Answer for Question 19:

Please refer to below image for billing amount calculation.

  • Cells highlighted in Grey are for data given in problem statement
  • Cells highlighted in Blue are for data computed from data given
  • Cells highlighted in Green are the outcomes of this question.


For Inbound handling, since total pallets intake is 300 which is less than 350, we will be considering 350 for billing perspective. Bill amount for inbound handling = 350 * 3.45 = $1207.50

Similarly, Bill amount for outbound handling = 350 * 5.52 = $ 1932.00

For Storage, since average daily storage is 120.714 pallets which is higher than minimum contractual obligation of 100 pallets, we will have to consider actual average daily storage i.e. 120.714 pallets

Hence, billing amount for storage = 120.714 * 13.8 = $ 1665.86

Total billing amount = $1207.50 + $ 1932.00 + $ 1665.86 = $ 4805.36